### v2

# April 4, 2025

Abstract. Televised debates between presidential candidates are often regarded as the exemplar of persuasive communication. Yet, recent evidence from Le Pennec and Pons (2023) indicates that they may not sway voters as strongly as popular belief suggests. We revisit their findings through the lens of the persuasion rate and introduce a robust framework that does not require exogenous treatment, parallel trends, or credible instruments. Instead, we leverage plausible monotonicity assumptions to partially identify the persuasion rate and related parameters. Our results reaffirm that the sharp upper bounds on the persuasive effects of TV debates remain modest.
Key Words: Persuasion rates; Monotonicity assumptions; Partial identification; Media; TV debates
JEL Classification Codes: C21, D72
1. Introduction
Economists have extensively documented empirical evidence demonstrating how persuasive efforts can affect the decisions and actions of consumers, voters, donors, and investors (see DellaVigna and Gentzkow, 2010, for a survey of the early literature). Indeed, the art of persuasion has captivated observers for centuries, from ancient rhetoricians to modern communication experts. In today’s media-driven world, few events draw more attention than a live TV debate between U.S. presidential candidates, often billed as a make-or-break moment that can sway millions of undecided voters. However, recent evidence from Le Pennec and Pons (2023, LP hereafter) challenges this common belief. Their large-scale study examines 56 TV debates across 31 elections in seven countries, and they find no discernible impact of those debates on vote choices, beliefs, or policy preferences.
In this paper, we revisit LP’s findings to reassess the effectiveness of TV debates from a different, robust perspective. Our framework focuses on the persuasion rate and related parameters, yet does not rely on linearity, additive separability, exogenous treatments, parallel trends, or credible instruments. Instead, it draws on a pair of monotonicity assumptions in the spirit of Manski and Pepper (2000). Because the sharp lower bounds are always trivially zero, we concentrate on the sharp upper bounds on the persuasion rate and other measures. Our results reaffirm LP’s conclusion in that these upper bounds remain small, implying that the overall effects of TV debates appear limited.
The persuasion rate, first introduced by DellaVigna and Kaplan (2007) to evaluate the efficacy of persuasive efforts, has been extensively used to make a comparison across diverse settings. Building upon this concept, Jun and Lee (2023) formulate the persuasion rate as a causal parameter that quantifies how a persuasive message affects recipients’ behavior, and they conduct a formal identification analysis based on exogenous treatment assignments or credible instruments. Extending this line of research, Jun and Lee (2024b) propose the average persuasion rate on the treated (APRT) and its reverse version (R-APRT) as additional causal parameters of interest, and they develop a difference-in-differences (DID) framework to identify, estimate, and conduct inference for them. Additional research on persuasion rates includes an analysis of persuasion types among compliers (Yu, 2025), identification of the average persuasion rate under sample selection (Possebom and Riva, 2025), an extension of the persuasion effect to continuous decisions (Kaji and Cao, 2025), and covariate-assisted bounds on the persuasion rate without monotonicity assumptions (Ji et al., 2024).
In this paper, we explore implications of two monotonicity assumptions, i.e., monotone treatment response (Manski, 1997, MTR) and monotone treatment selection (Manski and Pepper, 2000, MTS). The MTR assumption states that the effect of the informational treatment of interest is directional, and it has also been employed by Jun and Lee (2023, 2024b) in the context of the persuasion rates. In our application involving TV debates, it requires that uncertainty in vote intentions be lower after a debate takes place than before it does. Since debates provide additional information that should be helpful or at least not harmful, MTR appears plausible. For identifying the persuasion rate, MTR links the joint probability of two potential outcomes with the average treatment effect. The MTS assumption is new in studying the persuasion rate. In our context, it requires that the probability of vote choice consistency (voting exactly as intended at the time of the pre-election survey) not decrease as the survey date approaches election day, which seems natural.
Since the foundational work of Manski (1997) and Manski and Pepper (2000), employing MTR, MTS, and other types of monotonicity assumptions has been fruitful in the literature on partial identification. To sample a few, see Kreider and Pepper (2007), Vytlacil and Yildiz (2007), Bhattacharya et al. (2012), Kreider et al. (2012), Okumura and Usui (2014), Kim et al. (2018), Machado et al. (2019), and Jun and Lee (2024a) among others. Our paper is another example that shows the usefulness of the MTR and MTS assumptions.
It is worth noting that using MTR and MTS does not alter the estimands one would compute under MTR and exogenous treatments. What changes is how we interpret these estimands: the population quantities that point-identify the persuasion rate and its reverse version under MTR with exogenous treatments become the sharp upper bounds on these same measures when the exogenous treatment assumption is replaced with MTS. We illustrate these findings by reanalyzing LP’s data.
This robust interpretation of our estimands can be extended further. Both the persuasion rate and its reverse version are defined by conditioning on a counterfactual event. However, Pearl (1999, 2009, 2015) has advocated conditioning on an observable event rather than a counterfactual one. We show that this conceptual difference is inconsequential in our approach: our estimands serve as sharp upper bounds not only on the persuasion rate and its reverse version, but also on the probabilities of causation proposed by Pearl, when we solely rely on the MTR and MTS assumptions for identification.
Our empirical findings suggest that watching TV debates has, at most, a modest impact on vote choice consistency: the estimated average treatment effect is no higher than 0.7%, while the persuasion rate is about 3% at most. This implies that relatively few voters switch from an inconsistent to a consistent vote choice by watching debates. Our results align with those of LP, even though our methodology is quite different. We also remark that the upper bound on the persuasion rate is larger than that on the average treatment effect because, under the MTR assumption, the persuasion rate is derived by rescaling the average treatment effect. The rescaling factor specifically avoids “preaching to the converted” by conditioning on relevant events. However, even by this alternative measure, the effect of TV debates remains limited.
The rest of the paper is organized as follows. In section 2, we introduce the causal persuasion rate and related parameters, and discuss the MTR assumption, which will be used throughout. Section 3 focuses on identification, highlighting that the estimands that point-identify the parameters of interest under exogenous treatments become sharp upper bounds on those parameters when the MTS assumption is imposed. In section 4, we consider Pearl (1999)’s probabilities of causation and show that the same sharp upper bounds apply under MTR and MTS. In section 5, we apply our methodology to revisit LP’s data. In section 6, we provide concluding remarks. Finally, in the appendix, we give the proofs of our main theoretical results.
2. Preliminaries
For $d\in\{0,1\}$, let $Y(d)$ denote a binary potential outcome, $D$ the observed binary treatment, and therefore, the observed outcome is $Y=DY(1)+(1-D)Y(0)$. Because our identification results do not rely on covariates, we focus on the case without them. Alternatively, one can view our assumptions and identification results as “implicitly conditional on covariates.” The causal parameters we consider are:
$\theta=\mathbb{P}\{Y(1)=1\mid Y(0)=0\}\ \ \text{ and }\ \ \theta^{(r)}=\mathbb{P}\{Y(0)=0\mid Y(1)=1\}.$
The conditional probability $\theta$ is known as the average persuasion rate (APR), whereas $\theta^{(r)}$ is its reverse version (R-APR). Here, $\theta$ measures the proportion of individuals, among those who would not have taken the action of interest in the absence the treatment, who would have changed their behavior in the presence of the treatment. The reverse version $\theta^{(r)}$ focuses on individuals who would have taken the action of interest with the treatment, and it answers how many of them would not have done so without the treatment. See DellaVigna and Kaplan (2007), DellaVigna and Gentzkow (2010), and Jun and Lee (2023, 2024b) for more discussion on these and related parameters.
We start with making two assumptions.
Assumption A (Regularity).
For all $(y,d,d^{\prime})\in\{0,1\}^{3}$, $10<\mathbb{P}\{Y(d)=y,D=d^{\prime}\}<$ 1.
Assumption B (Monotone Treatment Response (MTR)).
We have $Y(1)\geq Y(0)$ with probability one.
Assumption A imposes standard regularity. Assumption B has been used for partial identification (e.g. Manski, 1997; Manski and Pepper, 2000; Jun and Lee, 2024a) or in the context of informational treatments (Jun and Lee, 2023, 2024b). In the latter case, it is often interpreted as an assumption that says, “Persuasive messages are directional, and there is no backlash.”
In the context of our application, assumption B can be justified by LP’s research design. In their study, respondents are surveyed twice, i.e., before and after the election: the first one is to capture their voting intentions, and the second to record their final choices. Our treatment variable $D$ is equal to $1$ if the first survey has occurred after a TV debate, and it is 00 otherwise. The outcome variable $Y$, termed vote choice consistency, is set to $1$ if a respondent’s pre-election vote intention matches their eventual vote, and 00 otherwise. In particular, $Y=0$ if a respondent’s initial response is “undecided.” Formally, let $Y(1)$ and $Y(0)$ represent vote choice consistency with and without a debate, respectively. Under the MTR assumption, we have $Y(1)\geq Y(0)$, meaning a debate cannot reduce vote consistency. For the sake of intuition, suppose that all cases of vote choice inconsistency correspond to undecided voters. Then, $Y(1)=0$ and $Y(0)=1$ means that this respondent would be an undecided voter in the presence of a debate but would know exactly how to vote without the debate. This scenario is implausible if watching a debate helps voters gain information and clarify their preferences. Therefore, we exclude that possibility. A similar reasoning applies to the other instances of $Y=0$.
Under assumptions A and B, the population consists of three (nonempty) groups: (i) never-persuadable (NP): $Y(1)=0,Y(0)=0$; (ii) already-persuaded (AP): $Y(1)=1,Y(0)=1$; and (iii) treatment-persuadable (TP): $Y(1)=1,Y(0)=0$. As we never observe the pair $\{Y(1),Y(0)\}$ jointly, we cannot assign a type to each observational unit; however, we can aim to identify the share of the three groups. Using the group probabilities, we can rewrite the parameters of interest as
$\theta=\frac{\mathbb{P}\{\mathrm{TP}\}}{\mathbb{P}\{\mathrm{TP}\}+\mathbb{P}\{\mathrm{NP}\}}\ \ \text{ and }\ \ \theta^{(r)}=\frac{\mathbb{P}\{\mathrm{TP}\}}{\mathbb{P}\{\mathrm{TP}\}+\mathbb{P}\{\mathrm{AP}\}}.$ (1)
We will now consider conditions under which we can point or partially identify the causal parameters.
3. Identification
We first consider the standard case of exogenous treatment as a benchmark.
Assumption C (Exogenous Treatment).
For $d\in\{0,1\}$, $Y(d)$ is independent of $D$.
Assumption A ensures that $\mathbb{P}(Y=1\mid D=1)>0$ and $\mathbb{P}(Y=1\mid D=0)<1$, and we define the following population quantities:
$\theta_{U}$ $=\frac{\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)}{\mathbb{P}(Y=0\mid D=0)},$
$\theta^{(r)}_{U}$ $=\frac{\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)}{\mathbb{P}(Y=1\mid D=1)}.$
Both $\theta_{U}$ and $\theta^{(r)}_{U}$ are directly identified from the data. We now state the following results.
Lemma 1 (Benchmark).
Under assumptions A, C and B, we have
$\mathbb{P}\{\mathrm{NP}\}=\mathbb{P}(Y=0\mid D=1)\quad\text{and}\quad\mathbb{P}\{\mathrm{AP}\}=\mathbb{P}(Y=1\mid D=0).$
In particular, $\theta$ and $\theta^{(r)}$ are identified by $\theta_{U}$ and $\theta^{(r)}_{U}$, respectively.
Proof.
The proof is straightforward and therefore it is omitted. ∎
Assumption C is popular, but it can be restrictive, as it precludes endogenous treatment assignment. Also, we are implicit here, but assumption C is often stated after conditioning on a sufficiently large vector of covariates, which can be too demanding in practice. To go beyond assumption C, we now consider an alternative assumption that allows endogenous treatment, namely monotone treatment selection.
Assumption D (Monotone Treatment Selection (MTS)).
For all $d\in\{0,1\}$, we have
$\mathbb{P}\{Y(d)=1\mid D=1\}\geq\mathbb{P}\{Y(d)=1\mid D=0\}.$
In LP’s study, the key identifying assumption is that, once potential confounders are accounted for, the date of the debate is uncorrelated with vote choice consistency. Therefore, roughly speaking, their event study design is based on (conditional) exogenous treatment assignment. However, controlling for all possible confounders in an observational study is inevitably challenging, even though LP makes a convincing case.
As an alternative, we employ the MTS assumption introduced by Manski and Pepper (2000). Formally, it requires that those in the treatment group be no less likely to “succeed” than those in the control group. Even if $D$ is a pure choice variable, assumption D stays valid as long as those who chose $D=1$ are not strictly inferior on average to those who opted out. In the context of LP, being surveyed later in the election cycle (i.e., $D=1$) naturally places respondents closer to the election day, and hence it makes vote consistency more likely, rendering MTS a plausible condition even without needing to control for covariates.
Theorem 1.
Suppose that assumptions A, B and D hold. Then, the sharp bounds on $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{AP}\}$, and $\mathbb{P}\{\mathrm{TP}\}$ are given by
$\mathbb{P}\{\mathrm{NP}\}\in\left[\mathbb{P}(Y=0\mid D=1),\mathbb{P}(Y=0)\right]\quad\text{and}\quad\mathbb{P}\{\mathrm{AP}\}\in\left[\mathbb{P}(Y=1\mid D=0),\mathbb{P}(Y=1)\right].$
In particular, the sharp identified intervals on $\theta,\theta^{(r)}$ are given by
$0\leq\theta\leq\theta_{U}\quad\text{ and }\quad 0\leq\theta^{(r)}\leq\theta^{(r)}_{U}.$
Proof.
See the appendix. ∎
Theorem 1 shows that, for both $\theta$ and $\theta^{(r)}$, the same population quantities $\theta_{U}$ and $\theta^{(r)}_{U}$ arise whether we assume assumption C or assumption D, although their interpretation changes. For example, under assumptions A, C and B, $\theta$ is point-identified by $\theta_{U}$, while that same $\theta_{U}$ becomes the sharp upper bound on $\theta$ under assumptions A, B and D. Consequently, a plug-in estimator of $\theta_{U}$ is consistent for $\theta$ under exogeneity, and it will be consistent for the sharp upper bound on $\theta$ under the weaker assumption of MTS.
Also, theorem 1 establishes bounds on the shares of the NP and AP groups. In particular, $\mathbb{P}(Y=0\mid D=1)$ and $\mathbb{P}(Y=1\mid D=0)$ are now interpreted as sharp lower bounds on $\mathbb{P}\{\mathrm{NP}\}$ and $\mathbb{P}\{\mathrm{AP}\}$, respectively. From the proof of theorem 1, we can see that assumptions B and D imply
$0\leq\mathbb{P}\{Y(1)=1\}-\mathbb{P}\{Y(0)=1\}\leq\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0),$
where the bounds are sharp. In other words, under assumptions B and D, the mean difference (the coefficient of $D$ when regressing $Y$ on $D$) serves as a sharp upper bound on the average treatment effect, and they coincide if the treatment is exogenous. Since $\mathbb{P}\{\mathrm{TP}\}=1-\mathbb{P}\{\mathrm{NP}\}-\mathbb{P}\{\mathrm{AP}\}=\mathbb{P}\{Y(1)=1\}-\mathbb{P}\{Y(0)=1\}$, the sharp bounds on $\mathbb{P}\{\mathrm{TP}\}$ are given by
$\mathbb{P}\{\mathrm{TP}\}\in\left[0,\ \mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)\right].$
We conclude this section with a brief discussion of estimation and inference. Plug-in estimators of $\theta_{U}$ and $\theta^{(r)}_{U}$ can be developed by regressing $Y$ on $D$ (including an intercept) and using the estimated coefficients. For inference, if assumption C is considered plausible, one can use the usual two-sided confidence intervals. Under the weaker assumption D, however, one-sided confidence intervals are more appropriate.
4. Bounding the Probabilities of Causation
In this section, we discuss a few related parameters that have been discussed in the causal inference literature. Both $\theta$ and $\theta^{(r)}$ are conditioned on a counterfactual outcome, which is not directly observed. Advocating the idea that we should consider probabilities of counterfactual events given the information that is observed, Pearl (1999) introduces the following parameters and calls them probabilities of causation:
$\mathrm{PS}$ =$=\mathbb{P}\{Y(1)=1\mid Y=0,D=0\},$
$\mathrm{PN}$ =$=\mathbb{P}\{Y(0)=0\mid Y=1,D=1\},$
$\mathrm{PNS}$ =$=\mathbb{P}\{Y(1)=1,Y(0)=0\},$
where PS, PN, and PNS, respectively, denote the probability of sufficiency, the probability of necessity, and the probability of necessity and sufficiency. For further discussion, see, e.g., Pearl (1999); Yamamoto (2012); Dawid et al. (2014); Dawid and Musio (2022); Zhang et al. (2024). These quantities can also be expressed as
$\mathrm{PS}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=0\}}{\mathbb{P}\{\mathrm{TP},D=0\}+\mathbb{P}\{\mathrm{NP},D=0\}},$
$\mathrm{PN}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=1\}}{\mathbb{P}\{\mathrm{TP},D=1\}+\mathbb{P}\{\mathrm{AP},D=1\}},$
$\mathrm{PNS}$ =$=\mathbb{P}\{\mathrm{TP}\}.$
Under assumptions A, C and B, we have that $\theta=\mathrm{PS}$, $\theta^{(r)}=\mathrm{PN}$, and $\mathrm{PNS}=\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0)$, as initially noted by Pearl (1999). More interestingly, under assumptions A, B and D, we have the following results.
Theorem 2.
Suppose that assumptions A, B and D hold. Then, the sharp identified intervals on $\mathrm{PS}$, $\mathrm{PN}$, and $\mathrm{PNS}$ are given by
$0\leq\mathrm{PS}\leq\theta_{U},\;0\leq\mathrm{PN}\leq\theta^{(r)}_{U},\;\text{ and }\;0\leq\mathrm{PNS}\leq\mathbb{P}(Y=1\mid D=1)-\mathbb{P}(Y=1\mid D=0).$
Proof.
See the appendix. ∎
Theorem 2 therefore shows that under assumptions A, B and D, the bounds on $\mathrm{PS}$ and $\mathrm{PN}$ coincide with those on $\theta$ and $\theta^{(r)}$, respectively. In other words, $\theta_{U}$ and $\theta^{(r)}_{U}$ also serve as the sharp upper bounds on $\mathrm{PS}$ and $\mathrm{PN}$, respectively. In addition, it confirms that $\mathrm{PNS}$ shares the same bounds as $\mathbb{P}\{\mathrm{TP}\}$, since both correspond to the average treatment effect.
5. Empirical Example
In this section, we revisit LP’s findings to reevaluate how effectively TV debates persuade (undecided) voters. The debates examined in LP took place between 5 and 44 days (on average 24 days) before the election, precisely when vote choice consistency tends to grow most swiftly. By employing an event-study design and capitalizing on the fact that debates occurred at varying times across different elections, LP control for the number of days preceding each election using daily fixed effects to distinguish the impact of the debates from general time trends. Interestingly, LP find no significant differences in how voters’ choices evolved before and after debate periods, rejecting any effect larger than $0.5$ percentage points. This null result holds for diverse voter types, different election contexts, and even tight, uncertain races.
In LP’s study, each observation corresponds to a respondent $\times$ debate $\times$ election, amounting to 331,000 total observations. The main specification of the regression model used in LP is (in their notation):
$Y_{it}^{d}=\sum_{k=-3}^{-1}\mu_{k}+\sum_{k=1}^{3}\mu_{k}+\mu_{4-}+\mu_{4+}+\sum_{t=-1}^{-60}\beta_{t}D_{t}+\alpha^{d}+W_{it}^{\prime}\lambda+u_{it}^{d},$ (2)
where $Y_{it}^{d}$ is the outcome for respondent $i$, surveyed $t$ days before the election linked to debate $d$. $D_{t}$’s, not to be confused with the treatment variable in our notation, are indicators to control for the number of days prior to the election, and the vector $W_{it}$ represents covariates such as day-of-the-week, postelectoral survey lag fixed effects, and sociodemographic characteristics. The terms $\mu_{k}(-3\leq k\leq 3)$ represent (the effects of) indicator variables for the number of days relative to the debate; $\mu_{4^{-}}$ and $\mu_{4_{+}}$ are indicators for being surveyed four or more days before or after the debate, respectively; and $\alpha^{d}$ are debate $\times$ election fixed effects.
The key coefficients of interest are $\mu_{1},\mu_{2}$, and $\mu_{3}$, which capture the effect of the debate one to three days afterward, relative to the omitted category $\mu_{0}$. They use the debate day as the reference because debates happen in the evening. They assume that any pre-trend appearing earlier than the fourth day preceding a debate or any effect emerging after the fourth day following a debate is adequately captured by the fixed effects $\mu_{4^{-}}$ and $\mu_{4^{+}}$.
Their identifying assumption is the one for the validity of the linear regression model in equation 2, which essentially requires that the date on which a debate occurred be unrelated to the outcome once conditioning on all other controls. Standard errors are clustered at the debate level.
In our empirical analysis, we take a different approach. Rather than controlling for a large set of variables as in (2), we restrict the sample to observations within a narrow time window of $[-3,+3]$ days relative to the debate. This localized approach allows us to focus on a period immediately surrounding the debate, where we define the pre-debate days as the control group ($D=0$) and the debate day along with the post-debate days as the treatment group ($D=1$).
We emphasize that we do not need to control for additional covariates, because, unlike the usual exogeneity assumption given covariates, the MTR and MTS assumptions do not require covariate adjustments for their validity, and they appear to be natural in this design as we discussed earlier. Specifically, MTR holds at the observational level, while MTS is likely to hold unconditionally. Since a debate provides information, “being undecided” is less likely in the presence of a debate, and hence MTR follows. Also, respondents surveyed later in the election cycle are naturally closer to the election day, which increases the likelihood of consistent voting behavior. Therefore, MTS follows as well. We remark that our approach does not rely on any functional form or exogeneity assumptions.
Table 1. Upper Bounds on the Persuasion Rates (in Percentage)
ATE APR ($\theta$) R-APR ($\theta^{(r)}$)
Est. SE UCB Est. SE UCB Est. SE UCB
All 0.68 0.56 1.60 3.37 2.76 7.91 0.84 0.69 1.98
Age $<$ 50 1.06 0.85 2.45 4.83 3.81 11.10 1.34 1.06 3.09
Age $\geq$ 50 0.40 0.58 1.36 2.13 3.09 7.22 0.49 0.71 1.66
U.S. 0.65 0.58 1.60 5.09 4.45 12.42 0.74 0.65 1.81
U.K. 0.61 0.90 2.08 3.69 5.43 12.63 0.72 1.06 2.46
Notes: The table presents the upper bound estimates (Est.), their standard errors (SE), and the 95% upper confidence bounds (UCB) for the average treatment effect (ATE), the average persuasion rate (APR) and its reverse version (R-APR), all expressed as percentages. Standard errors are clustered at the debate level, and the UCBs are computed using asymptotic normal approximations via the delta method.
Table 1 summarizes our empirical findings. In the first row, labeled “All,” we consider the full sample. The first three columns show the upper bound on the average treatment effect (ATE), its standard error, and the 95% upper confidence bound (UCB). The estimated upper bound of ATE is 0.7%, with a UCB of 1.6%. These results align closely with those in LP, although we rely on a notably different estimation method. Since the share of TP respondents is identical to ATE in our setting, these small values imply that very few voters would have consistent vote choices if only if they watched TV debates.
The middle three columns in Table 1 report the upper bounds on the average persuasion rate (APR, denoted by $\theta$), its standard error, and its 95% UCB, while the last three columns present analogous estimates for the reverse average persuasion rate (R-APR, $\theta^{(r)}$). Recall that APR and R-APR can be written as
$\textrm{APR}=\frac{\textrm{ATE}}{\textrm{share of TP}+\textrm{share of NP}},\quad\textrm{R-APR}=\frac{\textrm{ATE}}{\textrm{share of TP}+\textrm{share of AP}},$
where the denominators are less than one. Therefore, both APR and R-APR are larger than ATE, resulting in higher upper bounds. Here, our APR estimate exhibits a greater degree of amplification than R-APR, because the share of NP voters is smaller than that of AP voters. Economic interpretation of APR and R-APR in this application can be explained as follows. First, APR starts from those who would not have shown vote choice consistency without watching a debate, and it asks how many of them would have changed their behavior if they had watched a debate. In contrast, R-APR starts from those who would have shown vote choice consistency after watching a debate, and it asks how many of them would have been different if it were not for the debate. The former appears to be small, but the latter looks even smaller. Finally, we remark that the 95% confidence intervals are all one-sided because the lower bounds are trivially zero, and only the upper bounds need to be estimated (see theorem 1).
Subsequent rows of table 1 break down the analysis by subpopulations—respondents under 50, those 50 and older, and samples from the United States and the United Kingdom. Younger voters generally exhibit higher upper bounds than older voters, and the U.S. sample shows the largest upper bound on APR, suggesting a smaller share of APs in that subpopulation.
Overall, our findings indicate that TV debates exert, at best, a modest influence on vote choice consistency, echoing LP’s results. However, when focusing on APR (which excludes AP voters from the relevant population), the potential scope for debate effects is larger. Even so, the evidence remains suggestive because (i) the identified parameter is merely an upper bound on the true persuasive effect, and (ii) the standard errors associated with APR are relatively large.
6. Conclusions
We have revisited LP’s findings on the effects of TV debates by using a new framework involving the persuasion rate. Instead of relying on exogenous treatments or other strong identifying assumptions, we have leveraged two weak monotonicity assumptions (MTR and MTS) to derive sharp bounds on various measures of persuasive effects, including the persuasion rate and its reverse version. Our analysis has shown that the impact of TV debates on vote choice consistency remains modest, which is consistent with LP’s conclusion. In addition, we have highlighted that our bounds continue to serve as sharp bounds on Pearl’s probabilities of causation under the same monotonicity assumptions, further underscoring the robustness of our approach. Overall, our results suggest that while TV debates may provide helpful information to voters, their influence on electoral choices is limited.
Appendix
Proof of Theorem 1:
We first articulate the empirical contents of the MTR and MTS restrictions, after which we discuss identification of $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, $\mathbb{P}\{\mathrm{AP}\}$, $\theta$, and $\theta^{(r)}$.
Empirical Contents
We start from tabulating the joint distribution of $\bigl{(}Y(0),Y(1),D\bigr{)}$ as follows:
$Y(0),Y(1)$ $(0,0)$ $(0,1)$ $(1,0)$ $(1,1)$
$D=0$ $q_{1}$ $q_{2}$ 00 $q_{3}$
$D=1$ $q_{4}$ $q_{5}$ 00 $q_{6}$
Prob Restrictions $\sum_{j=1}^{6}q_{j}=1$.
$0\leq q_{j}\leq 1$ for $j=1,2,\cdots,6$.
Here, assumption B has already been imposed. Assumption D can be formulated by
$\frac{q_{6}}{q_{4}+q_{5}+q_{6}}$ $\geq\frac{q_{3}}{q_{1}+q_{2}+q_{3}}$ (3)
$\frac{q_{5}+q_{6}}{q_{4}+q_{5}+q_{6}}$ $\geq\frac{q_{2}+q_{3}}{q_{1}+q_{2}+q_{3}},$ (4)
where the denominators are not equal to zero by assumption A. Assumptions B and D do not impose any other restrictions on the $q_{j}$’s.
For $(y,d)\in\{0,1\}^{2}$, let $P_{yd}:=\mathbb{P}(Y=y,D=d)$, which are directly identified from the data. Then,
$\begin{split}q_{1}+q_{2}&=P_{00},\\ q_{4}&=P_{01},\end{split}\qquad\qquad\begin{split}q_{3}&=P_{10},\\ q_{5}+q_{6}&=1-P_{00}-P_{10}-P_{01},\end{split}$
where $\sum_{j=1}^{6}q_{j}=1$. Therefore, only two of the $q_{j}$’s are undetermined. Specifically, since $q_{6}=1-\sum_{j=1}^{5}q_{j}$, the last equation is redundant, and we can write
$q_{1}=P_{00}-q_{2},\quad q_{3}=P_{10},\quad q_{4}=P_{01},\quad q_{6}=P_{11}-q_{5}.$ (5)
That is, we have expressed all the $q_{j}$’s in terms of $q_{2}$ and $q_{5}$, where $0\leq q_{2}\leq P_{00}$ and $0\leq q_{5}\leq P_{11}$. Using equation 5, equations 3 and 4 can be rewritten as
$\frac{P_{11}-q_{5}}{P_{01}+P_{11}}$ $\geq\frac{P_{10}}{P_{00}+P_{10}},$ (6)
$\frac{P_{11}}{P_{01}+P_{11}}$ $\geq\frac{q_{2}+P_{10}}{P_{00}+P_{10}},$ (7)
from which we obtain
$0\leq q_{5}\leq P_{11}-\frac{P_{10}(P_{01}+P_{11})}{P_{00}+P_{10}},$ (8)
$0\leq q_{2}\leq P_{00}-\frac{P_{01}(P_{10}+P_{00})}{P_{01}+P_{11}}.$ (9)
Identification of $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, and $\mathbb{P}\{\mathrm{AP}\}$
Using the fact that
$\mathbb{P}\{\mathrm{NP}\}=q_{1}+q_{4},$$\;\mathbb{P}\{\mathrm{TP}\}=q_{2}+q_{5},\;\mathbb{P}\{\mathrm{AP}\}=q_{3}+q_{6},$
write $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, and $\mathbb{P}\{\mathrm{AP}\}$ in terms of $q_{2}$ and $q_{5}$ as before, so that
$\mathbb{P}\{\mathrm{NP}\}=P_{00}+P_{01}-q_{2},$$\;\mathbb{P}\{\mathrm{TP}\}=q_{2}+q_{5},\;\mathbb{P}\{\mathrm{AP}\}=P_{10}+P_{11}-q_{5},$
which is from equation 5. Recall that the bounds on $q_{2}$ and $q_{5}$ are a rectangle (see equations 8 and 9). Thus, the sharp bounds on $\mathbb{P}\{\mathrm{NP}\}$, $\mathbb{P}\{\mathrm{TP}\}$, and $\mathbb{P}\{\mathrm{AP}\}$ are
$\mathbb{P}\{\mathrm{NP}\}$ $\in\left[\mathbb{P}(Y=0\mid D=1),\mathbb{P}(Y=0)\right],$
$\mathbb{P}\{\mathrm{AP}\}$ $\in\left[\mathbb{P}(Y=1\mid D=0),\mathbb{P}(Y=1)\right],$
while $\mathbb{P}\{\mathrm{TP}\}=1-\mathbb{P}\{\mathrm{NP}\}-\mathbb{P}\{\mathrm{AP}\}$. ∎
Identification of $\theta$ and $\theta^{(r)}$
Now, note that
$\theta$ $=\frac{(q_{2}+q_{3}+q_{5}+q_{6})-(q_{3}+q_{6})}{1-(q_{3}+q_{6})}=\frac{q_{2}+q_{5}}{1-P_{10}-P_{11}+q_{5}},$ (10)
$\theta^{(r)}$ $=\frac{(q_{2}+q_{3}+q_{5}+q_{6})-(q_{3}+q_{6})}{q_{2}+q_{3}+q_{5}+q_{6}}=\frac{q_{2}+q_{5}}{q_{2}+P_{10}+P_{11}}.$ (11)
Therefore, we can obtain the upper bounds on $\theta$ and $\theta^{(r)}$ by maximizing equations 10 and 11, respectively, subject to equations 8 and 9; the lower bounds are similar.
First, consider $\theta$. The sharp lower bound on $\theta$ can be obtained by
[…]
Since $Q_{\theta}(\cdot,q_{5})$ is monotonic, it suffices to solve
[…]
from which it follows that the lower bound on $\theta$ is $Q_{\theta}(0,0)=0$. Similarly, for the upper bound on $\theta$, we consider
[…]
for which it suffices to consider
[…]
Therefore, it follows that the upper bound on $\theta$ is given by
$Q_{\theta}\left(P_{00}-\frac{P_{01}(P_{10}+P_{00})}{P_{01}+P_{11}},\ P_{11}-\frac{P_{10}(P_{01}+P_{11})}{P_{00}+P_{10}}\right).$
Let $Q_{d}:=P_{0d}+P_{1d}$, and simplifying this leads to
$\frac{P_{00}-\frac{P_{01}Q_{0}}{Q_{1}}+P_{11}-\frac{P_{10}Q_{1}}{Q_{0}}}{1-P_{10}-\frac{P_{10}Q_{1}}{Q_{0}}}$ $=\frac{Q_{0}P_{00}-\frac{P_{01}Q_{0}^{2}}{Q_{1}}+Q_{0}P_{11}-P_{10}Q_{1}}{Q_{0}-P_{10}}$
$=\frac{Q_{0}P_{00}-\frac{P_{01}Q_{0}^{2}}{Q_{1}}+Q_{0}P_{11}+P_{10}Q_{0}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{1}{Q_{1}}\left(Q_{0}Q_{1}P_{00}-P_{01}Q_{0}^{2}+Q_{0}Q_{1}P_{11}+Q_{0}Q_{1}P_{10}\right)-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{1}{Q_{1}}\left\{Q_{0}Q_{1}(1-P_{01})-P_{01}Q_{0}^{2}\right\}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{Q_{0}}{Q_{1}}(Q_{1}-P_{01})-P_{10}}{Q_{0}-P_{10}}$
$=\frac{\frac{Q_{0}}{Q_{1}}P_{11}-P_{10}}{Q_{0}-P_{10}}$
$=\frac{P_{11}/Q_{1}-P_{10}/Q_{0}}{1-P_{10}/Q_{0}}.$
Sharpness follows from continuity.
Now, for the upper bound on $\theta^{(r)}$, we solve
[…]
By the same reasoning as before, the maximum value is given by
$Q_{\theta}^{(r)}\left(P_{00}-\frac{P_{01}(P_{10}+P_{00})}{P_{01}+P_{11}},\ P_{11}-\frac{P_{10}(P_{01}+P_{11})}{P_{00}+P_{10}}\right)=\frac{P_{11}/Q_{1}-P_{10}/Q_{0}}{1-P_{01}/Q_{1}}.$
The lower bound on $\theta^{(r)}$ is easy, and it will be omitted.
Proof of Theorem 2:
First note that
$\mathrm{PS}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=0\}}{\mathbb{P}\{\mathrm{TP},D=0\}+\mathbb{P}\{\mathrm{NP},D=0\}},$
$\mathrm{PN}$ =$=\frac{\mathbb{P}\{\mathrm{TP},D=1\}}{\mathbb{P}\{\mathrm{TP},D=1\}+\mathbb{P}\{\mathrm{AP},D=1\}}.$
Using $q_{j}$’s defined in the previous proof, write
$\mathrm{PS}$ $=\frac{q_{2}}{q_{1}+q_{2}}=\frac{q_{2}}{P_{00}},$
$\mathrm{PN}$ $=\frac{q_{5}}{q_{5}+q_{6}}=\frac{q_{5}}{P_{11}}.$
Then, the sharp bounds on $\mathrm{PS}$ and $\mathrm{PN}$ are directly obtained from the bounds on $q_{2}$ and $q_{5}$, which are already given in equations 8 and 9. Finally, the sharp bounds on $\mathrm{PNS}$ are immediately obtained from theorem 2 because $\mathrm{PNS}=\mathbb{P}\{\mathrm{TP}\}$. ∎